HETAV PATEL

Grade 11,

difrentiate x(1+y)½ + y(1+x) ½ the answer get very soon lets see who will differentiate first

difrentiate x(1+y)½ + y(1+x)½                  
    the answer get  very soon lets see who will differentiate first                                                                  

Grade:11

1 Answers

Arun
5 years ago
Dear student
 

x√ (1+y) = -y √(1+x)
Sqare both sides
x^2(1+y) = (y^2)(1+x)
x^2 + y(x^2) = y^2 + x(y^2)

(x-y)(x+y) + xy(x-y) = 0
(x-y)(x+y+xy) = 0

x can't be equal to y , therefore
x+y+xy = 0
y = -x / (1+x)
Differentiate eqn w.r.t x

dy/dx = -1/(1+x)^2